# HAPNet: Toward Superior RGB-Thermal Scene Parsing via Hybrid, Asymmetric, and Progressive Heterogeneous Feature Fusion

### v2

@@ Abstract
Data-fusion networks have shown significant promise for RGB-thermal scene parsing. However, the majority of existing studies have relied on symmetric duplex encoders for heterogeneous feature extraction and fusion, paying inadequate attention to the inherent differences between RGB and thermal modalities. Recent progress in vision foundation models (VFMs) trained through self-supervision on vast amounts of unlabeled data has proven their ability to extract informative, general-purpose features. However, this potential has yet to be fully leveraged in the domain. In this study, we take one step toward this new research area by exploring a feasible strategy to fully exploit VFM features for RGB-thermal scene parsing. Specifically, we delve deeper into the unique characteristics of RGB and thermal modalities, thereby designing a hybrid, asymmetric encoder that incorporates both a VFM and a convolutional neural network. This design allows for more effective extraction of complementary heterogeneous features, which are subsequently fused in a dual-path, progressive manner. Moreover, we introduce an auxiliary task to further enrich the local semantics of the fused features, thereby improving the overall performance of RGB-thermal scene parsing. Our proposed HAPNet, equipped with all these components, demonstrates superior performance compared to all other state-of-the-art RGB-thermal scene parsing networks, achieving top ranks across three widely used public RGB-thermal scene parsing datasets. We believe this new paradigm has opened up new opportunities for future developments in data-fusion scene parsing approaches.
Index Terms:
@@
I Introduction
I-A Background
Unity in diversity strengthens perception. RGB-thermal (often abbreviated as RGB-T) scene parsing has emerged as a crucial feature in autonomous vehicles and mobile robots [1]. RGB images capture visible light, while thermal images capture heat signatures [2]. The fusion of these two modalities of data, especially during the feature encoding stage, has been proven to dramatically enhance the reliability and robustness of scene parsing [3, 4, 5]. Nevertheless, current data-fusion approaches [6, 7, 8, 9, 10, 11] generally employ duplex [12, 13, 14] encoders (two identical pre-trained hierarchical backbone networks with independent trainable parameters) to indiscriminately extract heterogeneous features from the RGB-T data, limiting their ability to fully exploit the advantages of both data modalities [15]. Therefore, this study aims to explore a more ingenious and effective strategy for heterogeneous feature extraction and fusion, with a specific focus on leveraging VFMs, as exemplified by BEiT series [16, 17] and DINOv2 [18], which have garnered considerable attention in recent computer vision research endeavors.
I-B Existing Challenges and Motivation
@@
Additionally, while symmetric duplex encoders [20, 21, 10] are prevalently used, they may not be the optimal architectures for heterogeneous feature extraction. This limitation arises from the substantial inherent modality differences [22] between the RGB and thermal images, as well as the unique advantages provided by vision Transformers (ViTs) and convolutional neural networks (CNNs). RGB images, rich in global semantic cues [23], are well-suited for learning through ViTs, while both RGB and thermal images provide gradient-related details (local semantics) [9, 24], which can be more effectively learned through CNNs. Therefore, developing a novel asymmetric duplex encoder, with ViT and CNN employed separately in each branch, for RGB-T scene parsing, stands as an underexplored area of research deserving greater attention.
@@
Moreover, the heterogeneous feature fusion strategy is of considerable importance in RGB-T scene parsing. The simplistic and indiscriminate strategies used in prior arts [4, 25, 5, 21] often cause conflicting feature representations and erroneous scene parsing results [15]. Taking MFNet [5] as an instance, its feature fusion relies solely on the element-wise concatenation of the heterogeneous features extracted from the duplex encoders at their final stage, treating the features from both modalities equally in the decoder’s input. Such a hard-coded feature fusion strategy overlooks the inherent differences in RGB-T feature characteristics and their respective reliability [9], resulting in unsatisfactory performance, particularly under poor illumination conditions. Therefore, another motivation for this article is to design an effective strategy for heterogeneous feature fusion, especially when these features are extracted from different modalities using an asymmetric duplex encoder.
I-C Contributions
@@
To address the aforementioned limitations, we investigate the use of VFMs for RGB-T scene parsing, thereby unleashing the power of self-supervised pre-training on extensive unlabeled data. Given the inherent differences between RGB and thermal images, we first develop a cross-modal spatial prior descriptor (CSPD) based on a CNN to capture robust local spatial patterns from both RGB and thermal images. We then combine the CSPD with a VFM to construct a hybrid, asymmetric encoder, which fuses the heterogeneous features (global context extracted from RGB images and spatial prior captured from RGB-T data) using a progressive heterogeneous feature integrator (PHFI). This dual-path, progressive strategy enables deeper fusion of heterogeneous features. Additionally, we introduce an auxiliary task to implicitly enrich the local semantics of the fine-grained fused features, thereby further improving the overall performance of RGB-T scene parsing. Our proposed Hybrid, Asymmetric, and Progressive Network (HAPNet) achieves the top rankings across three public RGB-T scene parsing datasets.
@@
  • We investigate the potential of VFM for RGB-T scene parsing and propose an asymmetric encoder that enables VFMs to achieve effective heterogeneous feature fusion.
  • By revisiting the inherent characteristics of RGB and thermal modalities, we design a hybrid, asymmetric encoder to fully exploit the strengths of both modalities.
  • We propose an auxiliary task to further enrich the local semantics of the fine-grained fused features.
@@
I-D Outline
The remainder of this article is structured as follows: Sect. II reviews related works. Sect. III details our proposed HAPNet. Sect. IV compares HAPNet with other SoTA methods and presents the ablation study results. Finally, Sect. V concludes this article and discusses possible future work.
II Related Work
@@
II-A Single-Modal Scene Parsing
The advent of the fully convolutional network (FCN) [26] catalyzed a wave of research on CNN-based scene parsing networks, which typically utilize convolutions for pixel-wise classification and explore various modules and architectures to enhance overall performance. For instance, the DeepLab series [27, 28] leverages atrous spatial pyramid pooling module to capture multi-scale context information, while the feature pyramid network (FPN) [29] effectively fuses multi-scale features in a fully convolutional fashion for robust dense prediction. Nonetheless, CNNs suffer from inherent limitations, e.g., constrained receptive fields, which impede their effectiveness in comprehensively modeling global context.
@@ II-A Single-Modal Scene Parsing
II-B Data-Fusion Scene Parsing
The aforementioned single-modal networks often face challenges in handling complex scenes, especially under poor illumination and adverse weather conditions [2]. This is primarily because they rely solely on color and texture information present in RGB images [9]. Therefore, researchers have explored multi-modal/source data-fusion techniques (with duplex encoders) to overcome these limitations. MFNet [5] pioneers the use of duplex encoders for RGB-T scene parsing. It extracts heterogeneous features using two lightweight backbones and realizes feature fusion via element-wise concatenation, achieving a balance between speed and accuracy. Most subsequent studies generally employ element-wise summation to achieve heterogeneous feature fusion and focus more on encoder and decoder designs for improved performance. For example, RTFNet [4] introduces skip connections in the decoder to preserve clear boundaries and fine details. Similarly, FuseSeg [25] replaces the ResNet [12] backbone in RTFNet with the stronger DenseNet [37], enabling richer multi-scale feature extraction. Despite the improved performance achieved by these networks, the simplistic heterogeneous feature fusion strategies potentially limit their ability to fully exploit the complementary information present in RGB-T data.
To address this drawback, recent works [21, 9, 8, 10, 6, 20, 7] have resorted to more advanced, learnable feature fusion strategies. For instance, FEANet [21] and GMNet [9] introduce attention-based modules to dynamically select important heterogeneous features. Their subsequent research endeavor, CMX [8], employs average and max pooling operations along with the attention modules to recalibrate heterogeneous features at the spatial and channel dimensions. Similarly, CAINet [10] employs different attention modules to model global context and aggregate local semantics across various spatial scales of RGB-T feature maps. Moreover, the ABMDRNet series [6, 20] introduces novel modality-wise bridging-then-fusing frameworks, capable of reconstructing pseudo images of one modality using the features extracted from the other modality. This approach successfully reduces modality discrepancy and generates more discriminative fused features. Additionally, inspired by the masked autoencoder (MAE) [38], CRM-RGBTSeg [7] randomly masks fixed-size regions within one modality while simultaneously providing complementary regions from another modality to duplex encoders, thus preventing over-reliance on either modality. In this article, we employ a novel asymmetric duplex encoder to capture both local spatial priors and global context from the RGB-T data. These heterogeneous features are progressively fused for more robust scene parsing.
III Methodology
@@ III-A Overall Workflow
Unlike previous works that utilize symmetric duplex encoder structures, we innovatively introduce an asymmetric, hybrid architecture (based on both a Transformer-based VFM and a lightweight CNN). It effectively leverages the complementary strengths of both modalities, thereby achieving deeper cross-modal feature fusion and further improving the overall performance. This is realized by dividing the ViT encoder into four blocks and inserting the PHFI to form four encoding stages. The RGB image $\boldsymbol{I}^{R}\in\mathbb{R}^{H\times W\times 3}$ and its corresponding thermal image $\boldsymbol{I}^{T}\in\mathbb{R}^{H\times W\times 3}$ are first fed into the CSPD, extracting the cross-modal spatial prior $\boldsymbol{F}^{P}_{1}\in\mathbb{R}^{(\sum_{i=2}^{4}{\frac{HW}{{S_{i}}^{2}})}\times D}$, which serves as one of the inputs to following stages, where $S_{i}=2^{i+1}$ ($i\in[2,4]\cap\mathbb{Z}$) denotes the corresponding stride number. Afterwards, $\boldsymbol{I}^{R}$ is tokenized to form the context feature $\boldsymbol{F}^{V}_{1}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$, which serves as the other input to following stages. $\boldsymbol{F}^{V}_{1}$ and $\boldsymbol{F}^{P}_{1}$ are then fused in a dual-path, progressive manner using the two key components of our developed PHFI: (1) the global-local context aggregator (GLCA) and (2) the complementary context generator (CCG). Meanwhile, fused features undergo global context encoding in four ViT blocks. This design enables the RGB features to retain fine-grained local semantics while effectively capturing global context through the powerful VFM. Eventually, the fused multi-scale features $\mathcal{F}^{F}=\{\boldsymbol{F}^{F}_{\frac{1}{S_{j}}}\in\mathbb{R}^{\frac{H}{S_{j}}\times\frac{W}{S_{j}}\times D}\}$, where $S_{j}=2^{i+1}$ ($j\in[1,4]\cap\mathbb{Z}$) from the four encoding stages are obtained and then fed into the subsequent mask classification decoder to generate semantic predictions $\boldsymbol{M}^{P}\in\mathbb{R}^{H\times W}$, which are also utilized by an auxiliary local semantics enhancement task during model training.
III-B Hybrid, Asymmetric Encoder
III-B1 Encoding Pipeline
@@
We evenly divide $\boldsymbol{I}^{R}$ into non-overlapping patches (resolution: $16\times 16$ pixels) and project them into $D$-dimensional tokens to form the RGB context feature $\boldsymbol{F}^{V}_{1}$. In the $i$-th encoding stage, the input context features $\boldsymbol{F}^{V}_{i}$ and spatial prior $\boldsymbol{F}^{P}_{i}$ are first fused in the GLCA to generate local semantics-enhanced context feature $\boldsymbol{\hat{F}}^{V}_{i}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$. Subsequently, $\boldsymbol{\hat{F}}^{V}_{i}$ undergoes context encoding in the ViT block, generating the output $\boldsymbol{F}^{V}_{i+1}$, which is then fed into the CCG to perform another fusion with $\boldsymbol{F}^{P}_{i}$. This step complements spatial prior $\boldsymbol{F}^{P}_{i}$ with updated global context from $\boldsymbol{F}^{V}_{i+1}$ to generate $\boldsymbol{F}^{P}_{i+1}$. By repeating such a dual-path feature fusion across four encoding stages, we obtain the fused spatial prior $\boldsymbol{F}^{P}_{5}$, which is then recovered to its original three resolutions, forming $\boldsymbol{F}^{F}_{\frac{1}{8}},\boldsymbol{F}^{F}_{\frac{1}{16}},\boldsymbol{F}^{F}_{\frac{1}{32}}\}$. A $2\times 2$ transposed convolution is employed to directly create the $\frac{1}{4}$-scale features $\boldsymbol{F}^{F}_{\frac{1}{4}}$ from $\boldsymbol{F}^{F}_{\frac{1}{8}}$. This design avoids the heavy computational overhead of the attention operations required for creating $\boldsymbol{F}^{F}_{\frac{1}{4}}$. Finally, these fused features across four scales constitute $\mathcal{F}^{F}$, which is compatible with current SoTA scene parsing network architectures [35, 39].
@@
Solely employing RGB images as the ViT input stems from our hypothesis that explicitly encoding thermal images using VFM may cause convergence issues and representation shift, as discussed in [40]. Furthermore, thermal data often has limited capability in capturing global context, as objects with similar geometries tend to have similar heat signatures. For example, distinguishing between a soccer ball and a basketball or between a cat and a dog is challenging in thermal images. We validate the effectiveness and superior performance of this hybrid, asymmetric network design through ablation studies in Sect. IV-E.
III-B2 Cross-modal Spatial Prior Descriptor

### v3

@@ Abstract
Data-fusion networks have shown significant promise for RGB-thermal scene parsing. However, the majority of existing studies have relied on symmetric duplex encoders for heterogeneous feature extraction and fusion, paying inadequate attention to the inherent differences between RGB and thermal modalities. Recent progress in vision foundation models (VFMs), which leverage self-supervised learning on large-scale unlabeled datasets, has exhibited superior capabilities in extracting informative, general-purpose features compared to supervised encoders. However, their potential has yet to be fully leveraged in the domain. In this study, we take one step toward this new research area by exploring a feasible strategy to fully exploit VFM features for RGB-thermal scene parsing. Specifically, we delve deeper into the unique characteristics of RGB and thermal modalities, thereby designing a hybrid, asymmetric encoder that incorporates both a VFM and a cross-modal spatial prior descriptor (CSPD), enabling enhanced extraction of complementary heterogeneous features. The extracted features undergo dual-path feature fusion through our proposed progressive heterogeneous feature integrators. Moreover, we introduce an auxiliary task to further enrich the local semantics of fused features, thereby improving the overall performance of RGB-thermal scene parsing. Our proposed HAPNet, incorporating all these components, delivers superior performance under challenging illumination conditions. Extensive experiments demonstrate that HAPNet outperforms all other state-of-the-art methods, with improvements of 0.1%, 1.0%, and 2.4% in mIoU on three public RGB-thermal scene parsing datasets: MFNet, PST900, and KP Day-Night, respectively. Additionally, our method exhibits exceptional generalizability for RGB-HHA scene parsing. We believe this new paradigm has opened up new opportunities for future developments in data-fusion scene parsing approaches.
keywords:
@@
journal: Biomimetic Intelligence and Robotics
[…]
[label1]organization=College of Electronic and Information Engineering, Tongji University, city=Shanghai, postcode=201804, country=China
[…]
[label2]organization=Department of Computer Science and Engineering, Hong Kong University of Science and Technology, city=Hong Kong, country=China
[…]
[label3]organization=MSU-BIT-SMBU Joint Research Center of Applied Mathematics, Shenzhen MSU-BIT University, city=Shenzhen, postcode=518172, country=China
[…]
[label8]organization=Qingdao Innovation and Development Base, Harbin Institute of Technology (Weihai), city=Qingdao, postcode=266109, country=China
[…]
[label4]organization=Department of Control Science and Engineering, Harbin Institute of Technology, city=Harbin, postcode=150001, country=China
[…]
[label6]organization=Harbin Institute of Technology at Weihai, city=Weihai, postcode=264200, country=China
[…]
[label7]organization=Suzhou Research Institute, Harbin Institute of Technology, city=Suzhou, postcode=215000, country=China
[…]
[label5]organization=Faculty of Computational Mathematics and Cybernetics, Lomonosov Moscow State University, city=Moscow, postcode=119991, country=Russia
1 Introduction
1.1 Background
Unity in diversity strengthens perception. RGB-thermal (often abbreviated as RGB-T) scene parsing has emerged as a crucial feature in autonomous vehicles and mobile robots [1]. RGB images capture visible light, while thermal images capture heat signatures. The fusion of these two modalities of data, especially during the feature encoding stage, has been proven to dramatically enhance the reliability and robustness of scene parsing [2, 3, 4]. Nevertheless, current data-fusion approaches [5, 6, 7, 8, 9, 10, 11, 12, 13] generally employ duplex [14, 15, 16, 17, 18] encoders (two identical pre-trained hierarchical backbone networks with independent trainable parameters) to indiscriminately extract heterogeneous features from the RGB-T data, limiting their ability to fully exploit the advantages of both data modalities [19]. Therefore, this study aims to explore a more ingenious and effective strategy for heterogeneous feature extraction and fusion, with a specific focus on leveraging VFMs, as exemplified by BEiT series [20, 21] and DINOv2 [22], which have garnered considerable attention in recent computer vision research endeavors.
1.2 Existing Challenges and Motivation
@@
Additionally, while symmetric duplex encoders [24, 25, 10] are prevalently used, they may not be the optimal architectures for heterogeneous feature extraction. This limitation arises from the substantial inherent modality differences [26] between the RGB and thermal images, as well as the unique advantages provided by vision Transformers (ViTs) and convolutional neural networks (CNNs). RGB images, rich in global semantic cues [27], are well-suited for learning through ViTs, while both RGB and thermal images provide gradient-related details (local semantics) [9], which can be more effectively learned through CNNs. Therefore, developing a novel asymmetric duplex encoder, with ViT and CNN employed separately in each branch, for RGB-T scene parsing, stands as an underexplored area of research deserving greater attention.
@@
Moreover, the heterogeneous feature fusion strategy is of considerable importance in RGB-T scene parsing [28]. The simplistic and indiscriminate strategies used in prior arts [3, 29, 4, 25] often cause conflicting feature representations and erroneous scene parsing results [19, 30]. Taking MFNet [4] as an instance, its feature fusion relies solely on the element-wise concatenation of the heterogeneous features extracted from the duplex encoders at their final stage, treating the features from both modalities equally in the decoder’s input. Such a hard-coded feature fusion strategy overlooks the inherent differences in RGB-T feature characteristics and their respective reliability [9], resulting in unsatisfactory performance, particularly under poor illumination conditions. Therefore, another motivation for this article is to design an effective strategy for heterogeneous feature fusion, especially when these features are extracted from different modalities using an asymmetric duplex encoder.
1.3 Contributions
@@
To address the aforementioned limitations, we propose the novel integration of VFMs into RGB-T scene parsing, thus unleashing the representational power derived from self-supervised pre-training on extensive unlabeled data. Given the inherent differences between RGB and thermal images, we first develop a cross-modal spatial prior descriptor (CSPD) based on a CNN to capture robust local spatial patterns from both RGB and thermal images. We then combine the CSPD with a VFM to construct a hybrid, asymmetric encoder, which fuses the heterogeneous features (global context extracted from RGB images and spatial prior captured from RGB-T data) using a progressive heterogeneous feature integrator (PHFI). To achieve this goal, two key components based on multi-scale deformable attention [31] are employed within each encoding stage in PHFIs. These components enable the effective integration of multi-scale spatial priors with single-scale global context. This dual-path, progressive strategy enables a deeper fusion of heterogeneous features. Additionally, inspired by the deep supervision approaches widely utilized in single-modal scene parsing tasks [32], we introduce an auxiliary task to implicitly enrich the local semantics of fine-grained fused features. This approach effectively reduces noisy and semantically irrelevant features, thereby further improving the overall performance of RGB-T scene parsing. Our proposed Hybrid, Asymmetric, and Progressive Network (HAPNet) outperforms all existing state-of-the-art (SoTA) RGB-T scene parsing methods that primarily leverage symmetric architectures. Specifically, HAPNet achieves improvements of 0.1%, 1.0%, and 2.4% in mean intersection over union (mIoU) on the MFNet [4], PST900 [1], and KP Day-Night [33] datasets, respectively. Furthermore, experiments on the NYU-Depth V2 [34] dataset demonstrate HAPNet’s generalizability for RGB-HHA scene parsing. These results demonstrate that HAPNet is particularly effective in handling a wide range of challenging scenarios, especially those characterized by poor illumination conditions or cluttered backgrounds.
@@
  • 1.
    We investigate the application of VFMs for RGB-T scene parsing, demonstrating that with appropriate strategies, methods in this domain can significantly benefit from VFMs.
  • 2.
    By revisiting the inherent characteristics of RGB and thermal modalities, we propose a hybrid, asymmetric encoder architecture as well as an asymmetric input design, which can fully exploit the complementary strengths of both modalities.
  • 3.
    We propose an auxiliary task to further enrich the local semantics of fine-grained fused features.
  • 4.
@@
1.4 Outline
The remainder of this article is structured as follows: Sect. 2 reviews related works. Sect. 3 details our proposed HAPNet. Sect. 4 compares HAPNet with other SoTA methods and presents the ablation study results. Finally, Sect. 5 concludes this article and discusses possible future work.
2 Related Work
@@
2.1 Single-Modal Scene Parsing
The advent of the fully convolutional network (FCN) [35] catalyzed a wave of research on CNN-based scene parsing networks, which typically utilize convolutions for pixel-wise classification and explore various modules and architectures to enhance overall performance. For instance, the DeepLab series [36, 37] leverages the atrous spatial pyramid pooling module to capture multi-scale context information, while the feature pyramid network (FPN) [38] effectively fuses multi-scale features in a fully convolutional fashion for robust dense prediction. Nonetheless, CNNs suffer from inherent limitations, e.g., constrained receptive fields, which impede their effectiveness in comprehensively modeling global context.
@@ 2.1 Single-Modal Scene Parsing
2.2 Data-Fusion Scene Parsing
The aforementioned single-modal networks often face challenges in handling complex scenes, especially under poor illumination and adverse weather conditions [28, 46]. This is primarily because they rely solely on color and texture information present in RGB images [9]. Therefore, researchers have explored multi-modal/source data-fusion techniques (with duplex encoders) to overcome these limitations. MFNet [4] pioneers the use of duplex encoders for RGB-T scene parsing. It extracts heterogeneous features using two lightweight backbones and realizes feature fusion via element-wise concatenation, achieving a balance between speed and accuracy. Most subsequent studies generally employ element-wise summation to achieve heterogeneous feature fusion and focus more on encoder and decoder designs for improved performance. For example, RTFNet [3] introduces skip connections in the decoder to preserve clear boundaries and fine details. Similarly, FuseSeg [29] replaces the ResNet [14] backbone in RTFNet with the stronger DenseNet [47], enabling richer multi-scale feature extraction. Despite their improved performance, these networks employ simplistic heterogeneous feature fusion strategies that limit their ability to fully exploit the complementary information in RGB-T data.
To address this drawback, recent works [25, 9, 8, 10, 5, 24, 7] have resorted to more advanced, learnable feature fusion strategies. For instance, FEANet [25] and GMNet [9] introduce attention-based modules to dynamically select important heterogeneous features. The subsequent research endeavor CMX [8], building upon the SegFormer [41] architecture, employs average and max pooling operations along with the attention modules to recalibrate heterogeneous features at the spatial and channel dimensions. Afterwards, its subsequent version CMNeXt [48] utilizes similar modules based on CNN and pooling operations for feature fusion, and further expands the architecture of CMX to accommodate an arbitrary number of modalities. Similarly, CAINet [10] employs different attention modules to model global context and aggregate local semantics across various spatial scales of RGB-T feature maps. Moreover, the ABMDRNet series [5, 24] introduces novel modality-wise bridging-then-fusing frameworks, capable of reconstructing pseudo images of one modality using the features extracted from the other modality. This approach successfully reduces modality discrepancy and generates more discriminative fused features. Additionally, inspired by the masked autoencoder (MAE) [49], CRM-RGBTSeg [7] randomly masks fixed-size regions within one modality while simultaneously providing complementary regions from another modality to duplex encoders, thus preventing over-reliance on either modality. In this article, we employ a novel asymmetric duplex encoder to capture both local spatial priors and global context from the RGB-T data. These heterogeneous features are progressively fused for more robust scene parsing.
3 Methodology
@@ 3.1 Overall Workflow
Unlike previous works that utilize symmetric duplex encoder structures, we innovatively introduce an asymmetric, hybrid architecture (based on both a Transformer-based VFM and a lightweight CNN). Compared to the indiscriminate feature extraction strategies utilized by symmetric duplex architectures, our HAPNet effectively leverages the complementary strengths of both modalities. It extracts rich global semantic cues using a VFM, while learning cross-modal local semantics with a CNN. Furthermore, HAPNet conducts effective global-local semantic integration within our proposed PHFIs, resulting in more distinguishable fused features. Consequently, our decoder can more effectively utilize the global context for accurate classification, while maintaining sensitivity to the local details of small-sized objects. This process starts by evenly dividing the ViT (with $L$ encoder layers) into four blocks, each of which contains $\frac{L}{4}$ encoder layers, and inserting PHFIs into each block to form four encoding stages. The RGB image $\boldsymbol{I}^{R}\in\mathbb{R}^{H\times W\times 3}$ and its corresponding thermal image $\boldsymbol{I}^{T}\in\mathbb{R}^{H\times W\times 3}$ are first fed into the CSPD, extracting the cross-modal spatial prior $\boldsymbol{F}^{P}_{1}\in\mathbb{R}^{(\sum_{i=2}^{4}{\frac{HW}{{S_{i}}^{2}})}\times D}$, which serves as one of the inputs to following stages, where $S_{i}=2^{i+1}$ ($i\in[2,4]\cap\mathbb{Z}$) denotes the corresponding stride number. Afterwards, $\boldsymbol{I}^{R}$ is tokenized to form the context feature $\boldsymbol{F}^{V}_{1}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$, which serves as the other input to following stages. $\boldsymbol{F}^{V}_{1}$ and $\boldsymbol{F}^{P}_{1}$ are then fused in a dual-path, progressive manner using the two key components of our developed PHFI: (1) the global-local context aggregator (GLCA) and (2) the complementary context generator (CCG). Meanwhile, fused features undergo global context encoding in four ViT blocks. This design enables the RGB features to retain fine-grained local semantics while effectively capturing global context through the powerful VFM. Eventually, the fused multi-scale features $\mathcal{F}^{F}=\{\boldsymbol{F}^{F}_{\frac{1}{S_{j}}}\in\mathbb{R}^{\frac{H}{S_{j}}\times\frac{W}{S_{j}}\times D}\}$, where $S_{j}=2^{i+1}$ ($j\in[1,4]\cap\mathbb{Z}$) from the four encoding stages are obtained and then fed into the subsequent mask classification decoder to generate semantic predictions $\boldsymbol{M}^{P}\in\mathbb{R}^{H\times W}$, which are also utilized by an auxiliary local semantics enhancement task during model training.
3.2 Hybrid, Asymmetric Encoder
3.2.1 Encoding Pipeline
@@
We evenly divide $\boldsymbol{I}^{R}$ into non-overlapping patches (resolution: $16\times 16$ pixels) and project them into $D$-dimensional tokens to form the RGB context feature $\boldsymbol{F}^{V}_{1}$. In the $i$-th encoding stage, the input context features $\boldsymbol{F}^{V}_{i}$ and spatial prior $\boldsymbol{F}^{P}_{i}$ are first fused in the GLCA to generate the local semantics-enhanced context feature $\boldsymbol{\hat{F}}^{V}_{i}\in\mathbb{R}^{\frac{HW}{16^{2}}\times D}$. Subsequently, $\boldsymbol{\hat{F}}^{V}_{i}$ undergoes context encoding in the ViT block, generating the output $\boldsymbol{F}^{V}_{i+1}$, which is then fed into the CCG to perform another fusion with $\boldsymbol{F}^{P}_{i}$. This step complements spatial prior $\boldsymbol{F}^{P}_{i}$ with updated global context from $\boldsymbol{F}^{V}_{i+1}$ to generate $\boldsymbol{F}^{P}_{i+1}$. By repeating such a dual-path feature fusion across four encoding stages, we obtain the fused spatial prior $\boldsymbol{F}^{P}_{5}$, which is then recovered to its original three resolutions, forming $\boldsymbol{F}^{F}_{\frac{1}{8}},\boldsymbol{F}^{F}_{\frac{1}{16}},\boldsymbol{F}^{F}_{\frac{1}{32}}\}$. A $2\times 2$ transposed convolution is employed to directly create the $\frac{1}{4}$-scale features $\boldsymbol{F}^{F}_{\frac{1}{4}}$ from $\boldsymbol{F}^{F}_{\frac{1}{8}}$. This design avoids the heavy computational overhead of the attention operations required for creating $\boldsymbol{F}^{F}_{\frac{1}{4}}$. Finally, these fused features across four scales constitute $\mathcal{F}^{F}$, which is compatible with current SoTA scene parsing network architectures [44, 50].
@@
Solely employing RGB images as the ViT input stems from our hypothesis that explicitly encoding thermal images using VFM may cause convergence issues and representation shift, as discussed in [51]. Furthermore, thermal data often has limited capability in capturing global context, as objects with similar geometries tend to have similar heat signatures. For example, distinguishing between a soccer ball and a basketball or between a cat and a dog is challenging in thermal images. We validate the effectiveness and superior performance of this hybrid, asymmetric network design through ablation studies in Sect. 4.5.
3.2.2 Cross-modal Spatial Prior Descriptor